OpenStudy (anonymous):

2. Which of the following has a graph that is wider than the graph of y = 3x² + 2?

OpenStudy (shayaan_mustafa):

r u there cutgirl?

OpenStudy (anonymous):

yea

OpenStudy (shayaan_mustafa):

so. give the following other graph as you said in question.

OpenStudy (anonymous):

y = 3x² + 3 y = 0.5x² + 1 y = –4x² – 1 y = 4x² + 1

OpenStudy (shayaan_mustafa):

y=0.5x2+1

OpenStudy (anonymous):

thanks

OpenStudy (shayaan_mustafa):

welcome.